PartnerObserving Log for 2007 Apr 18/19

Observers: A. Grazian, R. Gredel.

Night Assistant: D. Thompson, J. Hill

Telescope Operator: A. Ceranski

18:01 Turn on House Keeping Turn on Blue Channel

18:07 System ready. Temperature and pressure are OK. Launched sequence of 5 bias.

lbcb.20070419.011341.fits first bias lbcb.20070419.011556.fits last bias Lost one bias.

18:44 execute the 25bias.ob OB to take 25 bias frames

lbcb.20070419.014455.fits first bias lbcb.20070419.015825.fits last bias (lights on)

19:32 execute the dark900.ob OB to take a 900s dark image (lights off!!!) lbcb.20070419.023224.fits

Still too windy. Dome close.

19:53 Wind is going down. Ready to open the Dome.

20:05 Start with a Standard PG0918_UB (only point there). Bfastextra

20:31 Start with a Standard PG0918_UB Continue with PG0918_Vgr

Seeing 2.3arcsec; Photometric nights ?

21:02 Preset to fuobj8. Focus in B.

Start fuobj8_1_g lbcb.20070419.041045.fits lbcb.20070419.042820.fits

21:39 Preset to mpobj2_1_g and do image analysis.

21:54 Started OB mpobj2_1_g. lbcb.20070419.045519.fits Seeing 1.5 arcsec lbcb.20070419.050104.fits Guided on a galaxy lbcb.20070419.050656.fits lbcb.20070419.051257.fits Seeing is around 1 arcsec.

22:19 Started OB mpobj2_1_r. lbcb.20070419.051926.fits lbcb.20070419.052507.fits lbcb.20070419.053052.fits lbcb.20070419.053646.fits Seeing 1.2 arcsec.

22:48 Image analysis and then mpobj2_2_g. lbcb.20070419.054923.fits Seeing 1.1arcsec. lbcb.20070419.055505.fits lbcb.20070419.060051.fits lbcb.20070419.060645.fits Seeing 1.1arcsec.

LBC interface hangs up frequently. A warning message appears on the LOGs: "measured exposure time not available due to hardware timer failure, 10.057 sec extimated"

23:22 Rfastextra mpobj2_2_r

lbcb.20070419.060051.fits Seeing 1.1arcsec. lbcb.20070419.062859.fits Seeing 0.9 lbcb.20070419.063439.fits lbcb.20070419.064032.fits Seeing 0.9

23:52 Pointing M94. Bfastextra.ob; M94_NGC_4736_U (wind is high). lbcb.20070419.070751.fits Seeing worsen to 1.4arcsec. lbcb.20070419.071903.fits Seeing 1.6 Aborted observations.

00:26 Close the Dome for high wind.

01:42 The Dome is still close. JH started some "stress tests" for the shutter.

01:50 30 V exposures at horizon

02:30 30 R exposures at horizon

03:10 30 V exposures at zenith, 5 biases

03:45 30 R exposures at zenith, 5 biases

Shutter still working, wind still blowing.

-- JohnHill - 18-19 Apr 2007 ((Andrea Grazian as JH, and JH himself)
Topic revision: r5 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback